OpenStudy (anonymous):

10. What is y = -3/5x + 4 in standard form using integers?

jimthompson5910 (jim_thompson5910):

multiply every term by the LCD 5 to clear out the fractions to go from y = -3/5x + 4 to 5y = -3x + 20

jimthompson5910 (jim_thompson5910):

then get the equation into Ax + By = C form to get this 3x + 5y = 20

OpenStudy (anonymous):

thnxx!!